ann0818 - Announcement

Hubble status report

22 October 2008

Engineers at NASA continue to investigate the two anomalies which occurred during the reactivation of the NASA/ESA Hubble Space Telescope last week. Currently the focus is on assessing the ongoing risks for the activities needed to bring restore Hubble to full science capability prior to Servicing Mission 4.

In the last few days, engineers reviewed the data from both anomalies. On the Solar Blind Channel of the ACS the low voltage power supply reached its nominal 8 volt output just moments after the internal check. Engineers now need to determine if they can restart the power supply and whether there is any impact on the planned repair of ACS.

The cause of the Science Instrument Command and Data Handling (SI C&DH-B;) shutdown has been narrowed to an unspecified electrical event that triggered a reset of both Control Unit/Science Data Formatter-B and the NASA Standard Spacecraft Computer-1 (NSSC-1) Central Processing Module-B (CPM-B). This is the first time both these modules were activated on-orbit. Work to analyse the cause and determine potential workarounds is ongoing.

The main spacecraft remains in good health. The main subsystems are nominal and astrometry observations with the the Fine Guidance Sensors continue.

Image credit: NASA

Links

About Hubble instruments
About Hubble Servicing Mission 4

Contacts

Lars Lindberg Christensen
Hubble/ESA, Garching, Germany
Tel: +49-89-3200-6306
Cellular: +49-173-3872-621
E-mail: lars@eso.org

About the Announcement

Id:ann0818

Images

Hubble status report
Hubble status report

Also see our